# Unlearning Comparator: A Visual Analytics System for Comparative Evaluation of Machine Unlearning Methods

### v1

@@ V-A1 Retrained Model Behavior (Finding 1)
V-A2 Hyperparameter and Class-wise Trade-offs (Finding 2)
@@ V-A3 Misclassification Patterns (Finding 3)
After tuning the hyperparameters for each MU method, they became particularly interested in how confident each method was in its predictions for the forget class versus the retain classes (T3). To investigate this, they used the Prediction Matrix (Fig. 11). They found that most methods produced stable outputs (i.e., proportions closely matched their corresponding confidence) for the retain classes. However, the methods showed distinct differences in the forget class. While FT, GA, and SCRUB (Fig. 11B, D, and E) all yielded stable outputs, their prediction patterns diverged: FT and SCRUB were similar to the retrained model (Fig. 11A), whereas GA’s were not. Notably, RL and SalUn (Fig. 11C and F), both relying on random labeling, exhibited a clear mismatch between predicted proportions and confidence for the forget class (i.e., the brightness contrast between the two triangles in the first-row cells). This mismatch highlights an issue of poor confidence calibration [55]; such a model appears decisive due to high prediction proportions, while its low confidence values indicate internal uncertainty, leading to unreliable decisions in high-stakes domains such as medical diagnosis.
@@ VI Expert Feedback
Suggestions for Improvement. E1 and E3, both working in unlearning of classification models, offered suggestions on extending the system’s capabilities. E1 explained that it is difficult to monitor whether certain over-forgotten instances are being shifted to different classes, but the Embedding Space interactions made it simpler to track each sample’s changing predictions. Drawing from this, E1 suggested that visualizing embeddings from lower-level layers might offer useful insights. While we currently focus on the penultimate layer for efficient iterative workflow given the computational cost of processing high-dimensional lower-level layers, visualizing these layers could be a valuable option as computational efficiency improves. Also, we note that such 2D projections of high-dimensional data should be interpreted with caution, as projections can sometimes distort the original high-dimensional structure. In terms of dataset scalability, E3 noted that the current system is optimized for datasets with ten classes and suggested that supporting more classes would facilitate dataset expansion and tasks like sub-class unlearning. We acknowledge this scalability challenge as a promising future direction.
@@ VII Limitations and Future Work
Additionally, we gathered insights from both the researchers and existing literature regarding our system’s limitations and future research directions.
Supporting Different MU Tasks. Our interface currently focuses on class-wise unlearning, but researchers have asked about broader scenarios such as multi-class unlearning, instance-wise unlearning, or unlearning in generative models. In particular, unlearning for generative models requires handling knowledge distributed across outputs rather than discrete classes. Extending the interface to handle these would pose new design challenges.
Improving Privacy Evaluation. Quantifying privacy of MU methods remains challenging. Defining a privacy score and accurately estimating it in practice is still an open problem; this includes the distribution-level indistinguishability approach we adopted. Also, while our system focuses on black-box MIAs, we could consider scenarios where the attacker has access to the model weights (i.e., white-box attacks). Such an assumption would enable the evaluation of broader threats including model-inversion attacks [21] or attacks on the unlearning process itself through malicious unlearning requests [66]. Incorporating these scenarios may refine privacy evaluation and support more comprehensive assessments of MU methods.
Evaluation Without Retrained Models. Our current approach assumes access to a retrained model for comparison. In large-scale tasks such as LLM unlearning [16], however, retraining may be infeasible or prohibitively expensive. To address this concern, we could draw inspiration from the TOFU benchmark [67]. It asks GPT-4 to generate fake author profiles, fine-tunes a pre-trained LLM on them as the “original model,” applies unlearning to produce the “unlearned model,” and uses the pre-trained LLM (without fine-tuning) as the “retrained model.” Exploring design spaces that evaluate MU methods without a fully retrained model can be an important direction for future research.
@@ Acknowledgments
This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190421, Artificial Intelligence Graduate School Program (Sungkyunkwan University)) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00221186). This work was partly supported by Institute for Information & communication Technology Planning & evaluation (IITP) grants funded by the Korean government MSIT: (RS-2022-II220688 and RS-2021-II212068).

### v3

@@ V-A1 Retrained Model Behavior (Finding 1)
V-A2 Hyperparameter and Class-wise Trade-offs (Finding 2)
@@ V-A3 Misclassification Patterns (Finding 3)
After tuning the hyperparameters for each MU method, they became particularly interested in how confident each method was in its predictions for the forget class versus the retain classes (T3). To investigate this, they used the Prediction Matrix (Fig. 11). They found that most methods produced stable outputs (i.e., proportions closely matched their corresponding confidence) for the retain classes. However, the methods showed distinct differences in the forget class. While FT, GA, and SCRUB (Fig. 11B, D, and E) all yielded stable outputs, their prediction patterns diverged: FT and SCRUB were similar to the retrained model (Fig. 11A), whereas GA’s were not. Notably, RL and SalUn (Fig. 11C and F), both relying on random labeling, exhibited a clear mismatch between predicted proportions and confidence for the forget class (i.e., the brightness contrast between the two triangles in the first-row cells). This mismatch highlights an issue of poor confidence calibration [28]: high prediction proportions can mask low confidence, so these signals should be treated as diagnostic indicators and validated or calibrated before informing high-stakes decisions such as medical diagnosis.
@@ VI Expert Feedback
Suggestions for Improvement. E1 and E3, both working on unlearning for classification models, offered suggestions on extending the system’s capabilities. E1 explained that it is difficult to monitor whether certain over-forgotten instances are being shifted to different classes, but the Embedding Space interactions (e.g., linking corresponding instances) made it simpler to track each sample’s changing predictions. Drawing from this, E1 suggested that visualizing embeddings from lower-level layers might offer useful insights. While we currently focus on the penultimate layer for efficient iterative workflow given the computational cost of processing high-dimensional lower-level layers, visualizing these layers could be a valuable option as computational efficiency improves. Also, we note that such 2D projections of high-dimensional data should be interpreted with caution, as projections can sometimes distort the original high-dimensional structure. In terms of dataset scalability, E3 noted that the current system is optimized for datasets with ten classes and suggested that supporting more classes would facilitate dataset expansion and tasks like sub-class unlearning. We acknowledge this scalability challenge as a promising future direction.
@@ VII Limitations and Future Work
Additionally, we gathered insights from both the researchers and existing literature regarding our system’s limitations and avenues for future work.
Supporting Different MU Tasks. Our interface focuses on class-wise unlearning, but researchers have asked about broader scenarios such as multi-class unlearning, instance-wise unlearning, or sub-class unlearning. Specifically, for sub-class unlearning, our Prediction Matrix could be extended by adopting hierarchical confusion matrix designs as in Neo [26] to reveal inter-class misclassifications. Extending the interface to handle these would pose new design challenges.
Improving Privacy Evaluation. Quantifying privacy of MU methods remains challenging. Defining a privacy score and accurately estimating it in practice is still an open problem; this includes the distribution-level indistinguishability approach we adopted. Also, while our system focuses on black-box MIAs, we could consider scenarios where the attacker has access to the model weights (i.e., white-box attacks). Such an assumption would enable the evaluation of broader threats including model-inversion attacks [33] or attacks on the unlearning process itself through malicious unlearning requests [73]. Integrating these scenarios may refine privacy evaluation and support more comprehensive assessments of MU methods.
Extending to Foundation Model Unlearning. As MU research moves toward foundation models, extending our system will require redesign in several aspects. For data scalability, much larger datasets may benefit from complementing the instance-level Embedding Space view with embedding-based summaries, such as aggregated embeddings [31] or concept clustering [34], to support overview and navigation. For LLMs, supporting them would necessitate views for tracing unlearning-driven behavioral changes, such as token-level, sequence-oriented inspection [57] or attention-based behavior tracing [17], together with system optimizations to maintain interactivity under higher computational cost and latency. Moreover, our current system provides class-wise retrained models as gold-standard references; for foundation models, supplying such references in advance is often impractical at scale. As an alternative reference, training a retain-only baseline on MU benchmarks such as TOFU [42] could serve as a practical gold standard. Given that foundation model unlearning remains an active area with evolving evaluation conventions (e.g., OpenUnlearning [19]), incorporating such emerging practices can help align comparative MU analysis with real-world scenarios.
@@ Acknowledgments
This work was partly supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2023-00221186) and by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190421, Artificial Intelligence Graduate School Program (Sungkyunkwan University)). This work was partly supported by Institute for Information & communication Technology Planning & evaluation (IITP) grants funded by the Korean government MSIT: (RS-2022-II220688 and RS-2024-00437849).